OpenStudy (anonymous):

What are the next three terms in the sequence 16, 22, 28, 34, ...? 42, 48, 54 40, 46, 52 38, 44, 50

OpenStudy (ghazi):

40,46,52

OpenStudy (anonymous):

common differnce =6

Parth (parthkohli):

Add 6 to 34 Add 6 to 40 Add 6 to 46

OpenStudy (anonymous):

so just add 6 to the last number

Parth (parthkohli):

And keep messing around just like that. :)

OpenStudy (jiteshmeghwal9):

look carefully to the question , every number is adding six it it to gt the next term

OpenStudy (anonymous):

34 + 6=40 40+6 = 46 46+6 = 52 etc...

OpenStudy (anonymous):

thanks guysss

OpenStudy (anonymous):

so 40, 46, 52

OpenStudy (anonymous):

welcome........))

OpenStudy (jiteshmeghwal9):

first term:34+6=40 second term:40+6=46 third term:46+6=52